OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = −one halfx + 3 y = −one halfx + 6 y = 2x + 3

OpenStudy (anonymous):

@DebbieG @Luigi0210

OpenStudy (anonymous):

slope of required line is -1/2 so equation of line (y-3)=-1/2(x-6) so 2(y-3)=-(x-6) that is 2y=-x+6 that is y=-1/2x+6

OpenStudy (anonymous):

How do you get that ?

OpenStudy (anonymous):

Since, is perpendicular to the line y = 2x + 2. therefore the slope of the r3equired line is given as \[m=\frac{-1}{2}\] Since required line passes through the point (6, 3). therefore x1= 6 and y1= 3 Equation of line in slope point form is given as: (y-y1)=m(x-x1) (y-3)=-1/2(x-6) 2(y-3)=-1(x-6) 2y-6=-x+6 2y = -x+12 y= -1/2x+6 is the required eq of the line Thus y = −one halfx + 6 is the correcxt Answer

OpenStudy (anonymous):

@Sceenie

OpenStudy (anonymous):

Thank you .